OpenStudy (anonymous):

What are the explicit equation and domain for a geometric sequence with a first term of 3 and a second term of −9?

OpenStudy (anonymous):

an = 3(−12^)n − 1; all integers where n ≥ 1 an = 3(−12)^n − 1; all integers where n ≥ 0 an = 3(−3)^n − 1; all integers where n ≥ 1 an = 3(−3)^n − 1; all integers where n ≥ 0

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@ParthKohli @dpaInc @campbell_st @thomaster

OpenStudy (anonymous):

@robtobey @Luigi0210

OpenStudy (anonymous):

@ganeshie8